OpenStudy (anonymous):

help ?!? the last question for today. Circle A has a radius of 3 inches, and circle B has a radius of 7 inches. Prove that the two circles are similar.

OpenStudy (jdoe0001):

hmm... sounds like a odd question

OpenStudy (anonymous):

i know

OpenStudy (jdoe0001):

\(\bf \cfrac{\bigcirc Area1}{\bigcirc Area2}=\cfrac{\bigcirc component1^2}{\bigcirc component2^2}\implies \cfrac{\cancel{\pi} 3^2}{\cancel{\pi} 7^2}=\cfrac{(3)^2}{(7)^2}\) not sure if that helps, but would show the equal ratios

OpenStudy (anonymous):

@jdoe0001 thx alot